# Towards a Design Guideline for RPA Evaluation: A Survey of Large Language Model-Based Role-Playing Agents

### v1

@@ A Survey of Large Language Model-Based Role-Playing Agents
Wenyue Hua University of California, Santa Barbara                        Weimin Lyu Stony Brook University                        Toby Jia-Jun Li University of Notre Dame
Dakuo Wang thanks: Corresponding author: d.wang@northeastern.edu      Equal contribution.    Github repository: https://github.com/CRChenND/LLM_roleplay_agent_eval_survey Northeastern University
@@ 4.2 Task Attributes
Among them, simulated individuals and simulated society primarily use simulation as the research goal. Simulated individuals involve modeling specific individuals or groups, such as end-users Chen et al. (2024a), to study their behaviors and interactions in a controlled setting. Simulated Society focuses on social interactions, including cooperation Bouzekri et al. (2024), competition Wu et al. (2024), and communication Mishra et al. (2023), aiming to explore emergent social dynamics.

### v3

@@ A Survey of Large Language Model-Based Role-Playing Agents
Wenyue Hua University of California, Santa Barbara                        Weimin Lyu Stony Brook University                        Yanfang Ye University of Notre Dame
Toby Jia-Jun Li University of Notre Dame                        Dakuo Wang thanks: Corresponding author: d.wang@northeastern.edu      Equal contribution.   Github repository: https://github.com/CRChenND/LLM_roleplay_agent_eval_survey   Searchable webpage: https://agentsurvey.hailab.io/ Northeastern University
@@ 2 Related Work
Figure 2: Taxonomy of RPAs.
2.1 Taxonomy of RPAs
Existing literature Chen et al. (2024d); Tseng et al. (2024); Chen et al. (2024e); Mou et al. (2024a) classifies RPAs along two independent dimensions: Simulation Target and Simulation Scale. The Simulation Target dimension differentiates between agents that simulate specific individuals (e.g., historical figures, fictional characters, or individualized personas) and those that simulate group characteristics (e.g., artificial personas) Chen et al. (2024d); Tseng et al. (2024); Chen et al. (2024e). The Simulation Scale dimension categorizes agents by the complexity of their interactions, ranging from single-agent simulations with no social interaction to multi-agent systems that replicate structured or emergent societal behaviors Mou et al. (2024a).
To unify these perspectives, we introduce an integrated taxonomy for RPAs (Fig.2). The Simulation Target axis distinguishes between individual-focused and group-focused agents. Examples of individual-focused agents include digital twins, which model an individual’s decision-making process Rossetti et al. (2024), and personas, which emulate specific human-like characteristics Chen et al. (2024b). Group-focused agents include social simulacra, which model interactions between specific individuals within a group (e.g., the relationship dynamics in Detective Conan) Wu et al. (2024a), and synthetic societies, which replicate large-scale social structures and emergent group behaviors Park et al. (2023). The Simulation Scale axis differentiates between single-agent and multi-agent systems. Single-agent RPAs operate at an individual level, such as digital twins used for personalized recommendations or personas that generalize group characteristics for interaction. Multi-agent RPAs involve more complex interactions, with social simulacra capturing interpersonal dynamics within small, predefined groups, and synthetic societies modeling large-scale collective decision-making and societal structures.
2.2 Evaluation of RPAs
@@ 4.2 Task Attributes
Among them, simulated individuals and simulated society primarily use simulation as the research goal. Simulated individuals involve modeling specific individuals or groups, such as end-users Chen et al. (2024a), to study their behaviors and interactions in a controlled setting. Simulated Society focuses on social interactions, including cooperation Bouzekri et al. (2024), competition Wu et al. (2024b), and communication Mishra et al. (2023), aiming to explore emergent social dynamics.